OpenStudy (anonymous):

In a Chinese game show, 50% of the players are eliminated each day. If the show starts with 1,000 players, to the nearest integer, how many players will be present at the end of the 8th day of play? 1 4 14 41 what is the formula? help!

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow 1000(1 - {50 \over 100})^n }\]

OpenStudy (anonymous):

like is it the arithmetic sequences or geometric?

OpenStudy (anonymous):

Geometric.

Parth (parthkohli):

Geometric of course, as there is a common ratio.

OpenStudy (anonymous):

so is it the formula: or the geomertic series, or sequence?

OpenStudy (anonymous):

It is the formula of the 8th number of a geometric sequence, with common ratio 1/2, except that there is this phrase 'to the nearest integer'.

OpenStudy (anonymous):

is it the formula: an = a1(r)^(n-1) or is it a1 - a1(r)^n/1-r

OpenStudy (ash2326):

@sar12 did you understand?

OpenStudy (dumbcow):

its the first one to find 8th term in sequence

OpenStudy (anonymous):

understand what, i just need help with the formula.. because there is this n = a1(r)^(n-1) or this it a1 - a1(r)^n/1-r to do it and im confused which one....

OpenStudy (dumbcow):

the 2nd one deals with sum of geometric series

OpenStudy (anonymous):

so its an = a1(r)^(n-1)?

OpenStudy (anonymous):

a=1000 r=1/2

OpenStudy (anonymous):

yes and n = 8

OpenStudy (dumbcow):

actually n = 9 now that i look at question again....it says at END of 8th day

OpenStudy (anonymous):

just keep devide 8 times with a calculator and you get 3.90625 lol

OpenStudy (anonymous):

well which means when its finsihed not when its ninth day, right?

OpenStudy (anonymous):

Adumbcow, ur right i just did the nine for it and got 3.90625, which is 4.

OpenStudy (anonymous):

i mean @

OpenStudy (dumbcow):

:)